OpenStudy (anonymous):

Find the area of the triangle with base yards and height yards. Area of a triangle is equal to half the product of its base and height:

OpenStudy (anonymous):

use the theorem: \[a^2+b^2=c^2\] where c is he hyp, long side

OpenStudy (anonymous):

find the area of the triangle with the base \[2\frac{ 1 }{ 4 } \] and the height of \[\frac{ 4 }{ 7 }\] yards area of a triangle is equal to half the product

OpenStudy (cggurumanjunath):

area of triangle=0.5*base*height=0.5*(9/4)*(4/7)=9/14 or (0.64)yards^2

OpenStudy (cggurumanjunath):

is it clear or not ?

OpenStudy (anonymous):

the area is what yards squared

OpenStudy (cggurumanjunath):

0.64

OpenStudy (cggurumanjunath):

9/14

OpenStudy (cggurumanjunath):

9/14 or 0.64 both are same

OpenStudy (cggurumanjunath):